Not Sure taking a lab on a computer. Just want the answer (alreasy have it marked wrong) A 5lb bag of vegetables contains green beans (veg medealy) 5/9 of the bag is green beans. How many lbs of green beans are in the bag?
Answer by Mathtut(3670)

if 5/9 of the 5 pound bag are green beans then:
:
(5/9)*5=25/9=2 and 7/9 pounds of grean beans or 2.78 rounded to nearest 100th